# Finiteness criteria for the solutions of the sequence of semi-decomposable form equations and inequalities

### v1

@@ Abstract.
In this paper, we give some finiteness criteria for the solutions of the sequence of semi-decomposable form equations and inequalities, where the semi-decomposable form is factorized into a family of homogeneous polynomials with the distributive constant exceeding a certain number.
@@ 1. Introduction
Let $k$ be a number field and let $M_{k}$ be the set of places. Let $S$ be a subset of $M_{k}$ which contains all archimedean places of $k$. Denote by $\mathcal{O}_{S}$ and $O_{S}^{*}$ the set of all $S$-integers and the set of all $S$-units, respectively. Denote by $H_{S}({\bf x})$ the $S$-height of the point ${\bf x}\in k^{m+1}$ (see Section 2 for detail definitions).
Let $F(x_{1},\ldots,x_{m})$ be a decomposable form of degree $q$ which can be factorized into linear factors over a finite extension $k^{\prime}$ of $k$. An interesting problem in the diophantine approximation is to consider the solutions of the following decomposable form inequations (see [5, 10, 11, 12] for reference):
@@
Theorem A (see [3, Theorem 2]). Let $q,m$ be positive integers. Let $c,\lambda$ be real numbers with $c>0,\lambda<q-2m$, and $k^{\prime}$ a finite extension of $k$. For $n=1,2,\ldots$, let $F_{n}({\bf x})=F_{n}(x_{0},\ldots,x_{m})\in\mathcal{O}_{S}[{\bf x}]$ denote a decomposable form of degree $q$ which can be factorized into linear factors over $k^{\prime}$, and suppose that these factors are in general position for each $n$, which means if we consider the hyperplanes in $\mathbb{P}^{m}$ defined by these linear factors, they are in general position in $\mathbb{P}^{m}$. Then there does not exist an infinite sequence of $\mathcal{O}_{S}^{*}$-nonproportional ${\bf x}_{n}\in\mathcal{O}_{S}^{m+1},n=1,2,\ldots$, for which
$0<\prod_{v\in S}\|F_{n}({\bf x}_{n})\|_{v}\leqslant cH_{S}^{\lambda}({\bf x}_{n})$
and $h(F_{n})=o(h({\bf x}_{n}))\text{ as }n\rightarrow\infty.$
@@
In [9], by using the methods of Corvaja-Zannier [2] and Evertse-Ferretti [6, 7] and by introducing the notion of “distributive constant” of an arbitrary family of hypersurfaces with respect to a projective variety the author gave a general Schmidt’s subspace theorem. Motivated by these works, in this paper we will use the the notion of “distributive constant” to improve and generalize the above result to the case where each $F_{n}$ is a semi-decomposable form with a certain distributive constant. Here, a homogeneous polynomial $F(x_{0},\ldots,x_{m})\in k[x_{0},\ldots,x_{m}]$ is said to be semi-decomposable with the distributive constant $\Delta$ if $F$ can be factorized into a product of homogeneous polynomials $Q_{1},\ldots,Q_{q}$ over $\bar{k}$ with the distributive constant $\Delta$
In oder to state our result, we recall the following.
Definition 1.1.
Let $V$ be a projective subvariety of ${\mathbb{P}}^{m}$ defined over a number field $k$, and let $Q_{1},\ldots,Q_{q}$ be $q$ homogeneous polynomials in $\bar{k}[x_{0},\ldots,x_{m}]$. We define the distributive constant of the family $\mathcal{Q}=\{Q_{1},\ldots,Q_{q}\}$ with respect to $V(\bar{k})$ by
@@
$\Delta_{\mathcal{Q},V(\bar{k})}:={\mathrel{\mathop{{\max}}\limits_{{\emptyset\neq\Gamma\subset\{1,\ldots,q\}}}}}\dfrac{\sharp\Gamma}{\dim V(\bar{k})-\dim(\bigcap_{j\in\Gamma}Q_{j}^{*})\cap V(\bar{k})}.$
Throughout this paper, if $V={\mathbb{P}}^{m}$ then we just write $\Delta_{\mathcal{Q}}$ for $\Delta_{\mathcal{Q},{\mathbb{P}}^{m}(\bar{k})}$ and call it the distributive constant of $\mathcal{Q}$.
Here we note that $\dim\emptyset=-\infty$, by $\sharp\Gamma$ we denote the number of elements of the set $\Gamma$ and by $Q^{*}$ we denote the support of the homogeneous polynomial $Q$, ie.,
$Q^{*}=\{(x_{0}:\cdots:x_{m})\in{\mathbb{P}}^{m}(\bar{k})|Q(x_{0},\ldots,x_{m})=0\}.$
Remark 1.2.
(a) If $V^{\prime}$ is a subvariety of $V$ and $Q_{j}\not\subset V^{\prime}$ for all $j\in\{1,\ldots,q\}$ then
$\Delta_{\mathcal{Q},V^{\prime}(\bar{k})}\leqslant(\dim V(\bar{k})-\dim V^{\prime}(\bar{k})+1)\Delta_{\mathcal{Q},V(\bar{k})}.$
Indeed, since $V^{\prime}\subset V$ and $Q_{j}\not\subset V^{\prime}$ for all $j\in\{1,\ldots,q\}$, then for every nonempty subset $\Gamma$ of $\{1,\ldots,q\}$,
$(\dim V(\bar{k})-\dim V^{\prime}(\bar{k})+1)(\dim V^{\prime}(\bar{k})-\dim V^{\prime}(\bar{k})\cap\bigcap_{j\in\Gamma}Q_{j}^{*})$
$\geq\dim V(\bar{k})-\dim V^{\prime}(\bar{k})+\dim V^{\prime}(\bar{k})-\dim V^{\prime}(\bar{k})\cap\bigcap_{j\in\Gamma}Q_{j}^{*}$
$\geq\dim V(\bar{k})-\dim V(\bar{k})\cap\bigcap_{j\in\Gamma}Q_{j}^{*},$
i.e.,
$\frac{(\dim V(\bar{k})-\dim V^{\prime}(\bar{k})+1)\cdot\sharp\Gamma}{\dim V(\bar{k})-\dim V(\bar{k})\cap\bigcap_{j\in\Gamma}Q_{j}^{*}}\geq\frac{\sharp\Gamma}{\dim V^{\prime}(\bar{k})-\dim V^{\prime}(\bar{k})\cap\bigcap_{j\in\Gamma}Q_{j}^{*}}.$
This straightforwardly implies that $\Delta_{\mathcal{Q},V^{\prime}(\bar{k})}\leqslant(\dim V(\bar{k})-\dim V^{\prime}(\bar{k})+1)\Delta_{\mathcal{Q},V(\bar{k})}.$
(b) By the definition we also have $\Delta_{\mathcal{Q},V(\bar{k})}\geq 1$. If $Q_{1},\ldots,Q_{q}\ (q\geqslant\ell+1)$ are in $\ell-$subgeneral position with respect to $V$ then $\Delta_{\mathcal{Q},V(\bar{k})}\leq\ell-m+1$.
Our main result is stated as follows.
@@
Let $\ell,m,d$ be positive integers. Let $k^{\prime}$ be a finite extension of $k$ and $S\subset$ $M_{k}$ be a finite set containing all archimedean places. For $n=1,2,\ldots$, let $F_{n}({\bf x})=F_{n}(x_{0},\ldots,x_{m})\in\mathcal{O}_{S}[{\bf x}]$ be a sequence of semi-decomposable forms of degree $\ell$ with the distributive constant not exceeding a positive number $\Delta$ for each $n$, where $F_{n}=Q_{1,n}\cdots Q_{q,n}$ over $k^{\prime}$ with $\deg Q_{j,n}=d_{j,n}\ (1\leqslant j\leqslant q)$ such that $\max_{1\leqslant\jmath\leqslant q}d_{j,n}\leqslant d$ for each $n$ and $\Delta_{\{Q_{j,n}\}_{1\leq j\leq q}}\leq\Delta$ with $\ell>d\Delta\left(\frac{m}{2}+1\right)^{2}$. Let $c,\lambda$ be real numbers with $c>0,\lambda<\ell-d\Delta\left(\frac{m}{2}+1\right)^{2}$. Then, there does not exist an infinite sequence of $\mathcal{O}_{S}^{*}$-non-proportional $x_{n}\in\mathcal{O}_{S}^{m+1},n=1,2,\ldots$, for which
@@
(1.5) $h(F_{n})=o(h({\bf x}_{n}))\text{ as }n\rightarrow\infty$
@@
Let $k$ be a number field and $\{Q_{1},\ldots,Q_{q}\}$ a family of $q$ homogeneous polynomials in variables $(x_{1},\ldots,x_{m})$ with coefficients in $k$ and the distributive constant not exceeding a positive number $\Delta$. Let $d_{j}:=\deg Q_{j}=\ (1\leq j\leq q),d:=\max_{1\leqslant j\leqslant q}d_{j}$ and $F=Q_{1}\cdots Q_{q}$. Assume that $\deg F>d\Delta(\frac{m}{2}+1)^{2}$. Then, for every finite set $S$ of places of $k$ containing the archimedean places of $k$, for each positive number $\lambda<\deg F-d\Delta(\frac{m}{2}+1)^{2}$ and for each constant $c>0$, the inequality
@@
This corollary is an improvement and also is an generalization of Theorem 4.2 in [1].
@@ 2. Auxiliary result
Let $k$ be a number field. Denote by $M_{k}$ the set of places (equivalence classes of absolute values) of $k$ and by $M^{\infty}_{k}$ the set of archimedean places of $k$. For each $v\in M_{k}$, we choose the normalized absolute value $|\cdot|_{v}$ such that $|\cdot|_{v}=|\cdot|$ on $\mathbb{Q}$ (the standard absolute value) if $v$ is archimedean, and $|p|_{v}=p^{-1}$ if $v$ is non-archimedean and lies above the rational prime $p$. For each $v\in M_{k}$, denote by $k_{v}$ the completion of $k$ with respect to $v$ and set
@@
and the logarithmic $S$-height as $h_{S}({\bf x})=\log H_{S}({\bf x})$. If ${\bf x}\in\mathcal{O}_{S}^{m}\setminus\{0\}$, then $H_{S}({\bf x})\geq 1$ and $H_{S}(\alpha{\bf x})=H_{S}({\bf x})$ for all $\alpha\in\mathcal{O}_{S}^{*}$. It is clear that $H_{S}({\bf x})\geq h({\bf x})$ for any ${\bf x}\in\mathcal{O}_{S}^{m}\setminus\{0\}.$
Let =$\mathbb{N}=\{0,1,2,\ldots\}$ and for a positive integer $d$, we set
Let $Q=\sum_{I\in\mathcal{T}_{d}}a_{I}{\bf x}^{I}$ be a homogeneous polynomial of degree $d$ in $k[x_{0},\ldots,x_{m}]$, where ${\bf x}^{I}=x^{i_{0}}_{0}\ldots x^{i_{m}}_{m}$ for ${\bf x}=(x_{0},\ldots,x_{m})$ and $I=(i_{0},\ldots,i_{m})$. Define $\|Q\|_{v}=\max\{\|a_{I}\|_{v};I\in\mathcal{T}_{d}\}$. For each ${\bf x}=(x_{0},\ldots,x_{m})$, it is easy to see that
$\begin{cases}\|Q({\bf x})\|_{v}\leqslant\sharp\mathcal{T}_{d}\cdot\|Q\|_{v}\cdot\|{\bf x}\|_{v}&\text{ if }v\in M^{\infty}_{k},\\ \|Q({\bf x})\|_{v}\leqslant\|Q\|_{v}\cdot\|{\bf x}\|_{v}&\text{ if }v\in M_{k}\setminus M^{\infty}_{k}.\end{cases}$
The height of $Q$ is defined by

### v2

# Finiteness criteria for the solutions of a sequence of decomposable form inequalities

@@ Abstract.
In this paper, we give a finiteness criterion for the solutions of the sequence of semi-$q$-decomposable form equations and inequalities, where the semi-$q$-decomposable form is factorized into a family of $q$ nonconstant homogeneous polynomials with the distributive constant not exceeding a certain number.
@@ 1. Introduction
Let $k$ be a number field and let $M_{k}$ be the set of places. Let $S$ be a subset of $M_{k}$ which contains all archimedean places of $k$. Denote by $\mathcal{O}_{S}$ and $\mathcal{O}_{S}^{*}$ the set of all $S$-integers and the set of all $S$-units, respectively. Denote by $H_{S}({\bf x})$ the $S$-height of the point ${\bf x}\in k^{m+1}$ (see Section 2 for detailed definitions).
Let $F(x_{0},\ldots,x_{m})$ be a decomposable form of degree $q$ which can be factorized into linear factors over a finite extension $k^{\prime}$ of $k$. An interesting problem in Diophantine approximation is to consider the solutions of the following decomposable form inequations (see [6, 12, 13, 14] for reference):
@@
Theorem A (see [4, Theorem 2]). Let $q,m$ be positive integers. Let $c,\lambda$ be real numbers with $c>0,\lambda<q-2m$, and $k^{\prime}$ a finite extension of $k$. For $n=1,2,\ldots$, let $F_{n}({\bf x})=F_{n}(x_{0},\ldots,x_{m})\in\mathcal{O}_{S}[{\bf x}]$ denote a decomposable form of degree $q$ which can be factorized into $q$ linear factors over $k^{\prime}$, and suppose that these factors are in general position for each $n$, which means if we consider the hyperplanes in $\mathbb{P}^{m}$ defined by these linear factors, they are in general position in $\mathbb{P}^{m}$. Then there does not exist an infinite sequence of $\mathcal{O}_{S}^{*}$-nonproportional ${\bf x}_{n}\in\mathcal{O}_{S}^{m+1},n=1,2,\ldots$, for which
$0<\prod_{v\in S}\|F_{n}({\bf x}_{n})\|_{v}\leqslant cH_{S}^{\lambda}({\bf x}_{n})\ \text{ and }\ h(F_{n})=o(h({\bf x}_{n}))\text{ as }n\rightarrow\infty.$
In 2020, Ji, Yan and Yu generalized the above result of Györy and Ru by considering the case where each form $F_{n}$ is factorized into a product of homogeneous polynomials $Q_{1,n},\ldots,Q_{q,n}$ (see [9, Theorem 1.2]). To prove this result, they used the method of Yan and Yu in the paper [16], which involves constructing the Corvaja-Zannier filtration for the vector space of homogeneous polynomials vanishing on the image of the entire curves, in establishing the second main theorem for families of moving hypersurfaces in subgeneral position and nonconstant entire curves (see [16, Main Theorem]).
On the other hand, in 2022, by introducing the notion of “distributive constant” of an arbitrary family of hypersurfaces with respect to a projective variety the author gave a general Schmidt’s subspace theorem for fixed targets (see [11, Theorem 1.5]). Later on, in a very recent work [1], Cao, Thin and the author have given the notion of “distributive constant” for families of moving hypersurfaces indexed by an infinite index set and generalized our result in [11] to the case of sequence of targets (see [1, Theorem 1.6]).
@@
Motivated by these works, in this paper we will use the notion of “distributive constant” to improve and generalize Theorem A of Györy and Ru and [9, Theorem 1.2] of Ji-Yan-Yu to the case where each $F_{n}$ is factorized into a product of homogeneous polynomials $Q_{1,n},\ldots,Q_{q,n}$ and the family of moving hypersurfaces $\mathcal{Q}=\{Q_{1,n},\ldots,Q_{q,n}\}$ indexed by the set $\mathbb{N}$ of positive integers with a certain distributive constant. In order to state our result, we recall the following.
Let $\Lambda$ be an infinite index set. Let $\mathbb{N}=\{0,1,2,\ldots\}$ and for a positive integer $d$, we set
We define that a moving hypersurface $Q$ in ${\mathbb{P}}^{n}(k)$ of degree $d,$ indexed by $\Lambda$ is a collection of homogeneous polynomials $\{Q(\alpha)\}_{\alpha\in\Lambda}$ of the same degree $d$. Then, we can write
$Q=\sum_{I\in\mathcal{T}_{d}}a_{I}x^{I},$
where ${\bf x}^{I}=x^{i_{0}}_{0}\cdots x^{i_{m}}_{m}$ for ${\bf x}=(x_{0},\ldots,x_{m})$ and $I=(i_{0},\ldots,i_{m})$, and the $a_{I}$’s are maps (or functions) from $\Lambda$ to $k$ such that for every $\alpha\in\Lambda$, at least one $a_{I}(\alpha)$ is non-zero. We note that the support $Q(\alpha)^{*}=\{(x_{0}:\cdots:x_{n})\in{\mathbb{P}}^{n}(\bar{k})\ |\sum_{I\in\mathcal{T}_{d}}a_{I}(\alpha)x^{I}=0\}$ of $Q(\alpha)$ is a fixed hypersurface in ${\mathbb{P}}^{n}(\bar{k})$ for every $\alpha\in\Lambda$.
Definition B (see [1, Definition 1.5]). Let $V$ be a subvariety of $\mathbb{P}^{n}(k)$ of dimension $m$ and $\mathcal{Q}=\{Q_{1},\ldots,Q_{q}\}$ be a family of $q$ moving hypersurfaces of $\mathbb{P}^{n}(k)$ indexed by an infinite index set $\Lambda$. Assume that $V$ is not contained in $Q_{j}(\alpha)^{*}\;(j=1,\dots,q)$ for all, but finitely many $\alpha\in\Lambda$. We define the distributive constant of $\mathcal{Q}$ with respect to $V$ by the smallest number $\Delta_{\mathcal{Q},V}$ such that there exists an infinite index subset $A$ of $\Lambda$ with
@@
$\Delta_{\mathcal{Q},V}=\max_{\emptyset\neq\Gamma\subset\{1,\dots,q\},\alpha\in A}\dfrac{\#\Gamma}{m-\dim(\bigcap_{j\in\Gamma}Q_{j}(\alpha)^{*}\cap V(\overline{k}))}.$
Here we note that $\dim\emptyset=$-$\infty$, by $\sharp\Gamma$ we denote the number of elements of the set $\Gamma$. From the definition it follows that $\Delta_{\mathcal{Q},V}\geq 1$. Moreover, if $Q_{1},\ldots,Q_{q}\ (q\geqslant\ell+1)$ are in $\ell$-subgeneral position with respect to $V$ then $\Delta_{\mathcal{Q},V}\leq\ell-m+1$ (see [1]). In the case $V=\mathbb{P}^{n}(k)$, we just call $\Delta_{\mathcal{Q},\mathbb{P}^{n}(k)}$ the distributive constant of the family $\mathcal{Q}$.
Now, let $q$ be a positive integer and $F\left(x_{0},\ldots,x_{m}\right)$ a homogeneous polynomial in $k\left[x_{0},\ldots,x_{m}\right]\ (m\geq 1)$. We say that $F$ is a semi-$q$-decomposable form if $F$ can be factorized into a product of $q$ nonconstant homogeneous polynomials $Q_{1},\ldots,Q_{q}$ over $\bar{k}$. Our main result is stated as follows.
@@
Let $\ell,m,q$ be positive integers and $\Delta\geq 1$ a positive constant. Let $k^{\prime}$ be a finite extension of $k$ and $S\subset$ $M_{k}$ be a finite set containing all archimedean places. For $n=1,2,\ldots$, let $F_{n}({\bf x})=F_{n}(x_{0},\ldots,x_{m})\in\mathcal{O}_{S}[{\bf x}]$ be a sequence of semi-$q$-decomposable forms of degree $\ell$. Assume that $F_{n}=Q_{1,n}\cdots Q_{q,n}$ over $k^{\prime}$ with $\deg Q_{j,n}=d_{j},1\leqslant j\leqslant q$, for each $n$. Let $d=\max_{1\leqslant j\leqslant q}d_{j}$. Assume that $\ell>d\Delta\left(\frac{m}{2}+1\right)^{2}$ and the family of $q$ moving hypersurfaces $\{Q_{1,n},\ldots,Q_{q,n}\}$ (indexed by $\mathbb{N}$) has distributive constant not exceeding $\Delta$ for each $n$. Let $c,\lambda$ be real numbers with $c>0,\lambda<\ell-d\Delta\left(\frac{m}{2}+1\right)^{2}$. Then, there does not exist an infinite sequence of $\mathcal{O}_{S}^{*}$-non-proportional $x_{n}\in\mathcal{O}_{S}^{m+1},n=1,2,\ldots$, for which
@@
(1.3) $h(F_{n})=o(h({\bf x}_{n}))\text{ as }n\rightarrow\infty.$
@@
Let $k$ be a number field and $\{Q_{1},\ldots,Q_{q}\}$ a family of $q$ homogeneous polynomials in variables $(x_{1},\ldots,x_{m})$ with coefficients in $k$ and which has the distributive constant not exceeding a positive number $\Delta$. Let $d_{j}:=\deg Q_{j}\ (1\leqslant j\leqslant q)$ and $d:=\max_{1\leqslant j\leqslant q}d_{j}$ and $F=Q_{1}\cdots Q_{q}$. Assume that $\deg F>d\Delta(\frac{m}{2}+1)^{2}$. Then, for every finite set $S$ of places of $k$ containing the archimedean places of $k$, for each positive number $\lambda<\deg F-d\Delta(\frac{m}{2}+1)^{2}$ and for each constant $c>0$, the inequality
@@
This corollary is an improvement and also a generalization of Theorem 4.2 in [2].
@@ 2. Auxiliary result
Let $k$ be a number field. Let $M_{k}$ be the set of places of $k$ and $M^{\infty}_{k}$ be the set of Archimedean places. For each $v\in M_{k}$, we choose the normalized absolute value $|\cdot|_{v}$ such that $|\cdot|_{v}=|\cdot|$ on $\mathbb{Q}$ (the standard absolute value) if $v$ is archimedean, and $|p|_{v}=p^{-1}$ if $v$ is non-archimedean and lies above the rational prime $p$. For each $v\in M_{k}$, denote by $k_{v}$ the completion of $k$ with respect to $v$ and set
@@
and the logarithmic $S$-height as $h_{S}({\bf x})=\log H_{S}({\bf x})$. If ${\bf x}\in\mathcal{O}_{S}^{m}\setminus\{0\}$, then $H_{S}({\bf x})\geqslant 1$ and $H_{S}(\alpha{\bf x})=H_{S}({\bf x})$ for all $\alpha\in\mathcal{O}_{S}^{*}$. It is clear that $H_{S}({\bf x})\geqslant h({\bf x})$ for any ${\bf x}\in\mathcal{O}_{S}^{m}\setminus\{0\}.$
Let $Q=\sum_{I\in\mathcal{T}_{d}}a_{I}{\bf x}^{I}$ be a homogeneous polynomial of degree $d$ in $k[x_{0},\ldots,x_{m}]$. The support $Q^{*}$ of $Q$ is a fixed hypersurface in ${\mathbb{P}}^{n}(\bar{k})$ defined by
$Q^{*}=\{(x_{0}:\cdots:x_{n})\in{\mathbb{P}}^{n}(\bar{k})\ |\sum_{I\in\mathcal{T}_{d}}a_{I}x^{I}=0\},$
where $\bar{k}$ is the algebraic closure of the field $k$. The norm of $Q$ under $v\in S$ is defined by
$\|Q\|_{v}=\max\{\|a_{I}\|_{v};I\in\mathcal{T}_{d}\}.$
For each ${\bf x}=(x_{0},\ldots,x_{m})$, it is easy to see that
$\begin{cases}\|Q({\bf x})\|_{v}\leqslant\sharp\mathcal{T}_{d}\cdot\|Q\|_{v}\cdot\|{\bf x}\|^{d}_{v}&\text{ if }v\in M^{\infty}_{k},\\ \|Q({\bf x})\|_{v}\leqslant\|Q\|_{v}\cdot\|{\bf x}\|^{d}_{v}&\text{ if }v\in M_{k}\setminus M^{\infty}_{k},\end{cases}$
where $\sharp\mathcal{T}_{d}$ is the cardinalty of $\mathcal{T}_{d}$. The height of $Q$ is defined by